# VideoChat-R1: Enhancing Spatio-Temporal Perception via Reinforcement Fine-Tuning

### v2

@@ 4.1 Evaluation of VideoChat-R1
Method Charades-STA ActivityNet NExTGQA GoT VideoMME MVBench Peception Test mIoU R@0.5 R@0.7 mIoU R@0.5 R@0.7 mIoU acc Overlap R@0.5 Short-Avg Avg Val Baseline Qwen2.5-VL-7B 29.0 24.2 11.1 21.1 15.8 7.5 15.4 59.5 12.6 1.1 71.3 66.9 69.1 SFT on specific tasks +SFT w/ Charades-STA 46.3 63.9 45.0 25.3 20.6 30.2 16.7 7.9 - - N/A* N/A* N/A* +SFT w/ GoT - - - - - - - - 41.8 29.5 59.2 58.6 58.5 +SFT w/ NextQQA - - - - - - 28.2 64.8 - - 60.1 59.2 60.7 GRPO on various tasks VideoChat-R1 60.8 71.7 50.2 36.6 33.4 17.7 32.4 70.6 43.8 38.2 72.2 67.9 70.0
@@ 4.2 Ablation Studies and Discussions
Method GoT VideoMME Average overlap R@0.5 R@0.7 Short-Avg Qwen2.5-VL-7B 12.6 1.1 0 71.3 +SFT 41.8 29.5 3.9 59.2 +GRPO 42.5(+29.2) 30.6(+18.0) 3.9(+2.8) 71.4(+0.1)

### v3

@@ 4.1 Evaluation of VideoChat-R1
Method Charades-STA ActivityNet NExTGQA GoT VideoMME MVBench Peception Test mIoU R@0.5 R@0.7 mIoU R@0.5 R@0.7 mIoU acc Overlap R@0.5 Short-Avg Avg Val Baseline Qwen2.5-VL-7B 29.0 24.2 11.1 21.1 15.8 7.5 15.4 59.5 12.6 1.1 71.3 66.9 69.1 SFT on specific tasks +SFT w/ Charades-STA 46.3 45.0 25.3 20.6 16.7 7.9 - - - - N/A* N/A* N/A* +SFT w/ GoT - - - - - - - - 41.8 29.5 59.2 58.6 58.5 +SFT w/ NextQQA - - - - - - 28.2 64.8 - - 60.1 59.2 60.7 GRPO on various tasks VideoChat-R1 60.8 71.7 50.2 36.6 33.4 17.7 32.4 70.6 43.8 38.2 72.2 67.9 70.0 VideoChat-R1-thinking 59.9 70.6 47.2 35.5 33.3 16.7 36.1 69.2 43.3 33.9 74.2 66.2 69.6
@@ 4.2 Ablation Studies and Discussions
Method GoT VideoMME Average overlap R@0.5 R@0.7 Short-Avg Qwen2.5-VL-7B 12.6 1.1 0 71.3 +SFT 41.8 29.5 3.9 59.2 +GRPO 42.5(+29.2) 30.6(+29.5) 3.9(+3.9) 71.4(+0.1)